# A Manifold Perspective on the Statistical Generalization of Graph Neural Networks

### v2

@@ Appendix G Further references
Related to the manifold model we consider, some works construct graphs sampled from graphons, which can be seen as a random limit graph model. This line of work has concentrated on their convergence, stability and transferability properties [53, 54, 55]. Graphon is also used as a tool for pooling in GNNs in [56]. Albeit useful as a model, the graphon presents several limitations compared to the manifold model we use. Firstly, the graphon model assumes an infinite degree at every node [57], which is not the case in the manifold model. Secondly, the graphon presents little intuition on the underlying model, outside of the stochastic block model graphon, it is not trivial to visualize a graphon. On the other hand, manifolds are more interpretable, as can be the case on the sphere, or on other shapes (see Figure 2). Lastly, the manifold model is a more realistic assumption in the types of models that it can characterize.
@@
We consider an underlying manifold model from which we sample to obtain graphs, which are later the inputs to the GNN. Related to us, some works have used the sampled graphs the infer properties about the manifold itself. In particular, some works have set out to test the manifold assumption of the data [58], the manifold dimension [59] and show the complexity of doing so [60, 61]. In terms of prediction and classification using manifolds, and manifold data, several algorithms and methods have been proposed – notable the Isomap algorithm [62, 63, 64] and other manifold learning techniques [65]. In all, these techniques utilize the manifold assumption to infer properties of the manifold, which differs from our goal of providing a generalization bound for GNNs.
@@
Transferability of GNNs has been analyzed in many works by comparing the output differences of GNNs on different sizes of graphs when graphs converge to a limit model without statistical generalization analysis. In [15, 53, 54], the transferability of GNNs on graphs generated from graphon models is analyzed with difference bounds of GNN outputs proved. In [66], the authors show how increasing the size of the graph as the GNN learns, generalizes to the large-scale graph. In [67], the authors analyze the transferability of GNNs on graphs generated from a general topological space while the authors in [35] focus on graphs generated from manifolds. In [68], the authors propose a novel graphop operator as a limit model for both dense and sparse graphs with a transferability result proved. In [69] and [70], the authors study the transfer learning of GNNs with measurements of distance between graphs without a limit assumption. In [71], a transferable graph transformer is proposed and proved with empirical results.
@@ Appendix H Low Pass Filter Assumption
In the main results, we assume that the GNN and MNN are low-pass filters. We believe this is a mild assumption. A high frequency on the graph/manifold implies that there exist signals (eigenvectors/functions) that can have large variations in adjacent entries. This naturally generates instabilities, which are more difficult to learn. It is expected certain amount of homogeneity within the neighborhood of a node, translates into low-frequency signals. It has been shown in practice that several real-world physical models lead to low-pass filters such as opinion dynamics [72], econometrics [73], and graph signal processing [74].
Several other learning techniques rely on low-pass filters. Principal Component Analysis (PCA), can be seen as low-pass filters applied to the correlation matrix. Isomap, is another example of low pass filtering, but in this case applied to the Laplacian matrix. In all, we believe that the low pass filter assumption is not a hard assumption, given its empirical verification, and the fact that several other techniques that work in practice utilize it.

### v3

@@ Appendix G Further references
Different from the manifold model we are using, some research constructs graphs derived from graphons, which can be viewed as a random limit graph model. This research has focused on their convergence, stability, as well as transferability [53, 54, 55]. In [56], graphon is used as a pooling tool in GNNs. Despite its utility, the graphon presents several limitations compared to the manifold model we use. Firstly, the graphon model assumes an infinite degree at every node [57], which is not the case in the manifold model. Additionally, graphons offer limited insight into the underlying model; visualizing a graphon is challenging, except in the stochastic block model case. Manifolds, however, are more interpretable, especially when based on familiar shapes like spheres and 3D models (see Figure 2). Finally, the manifold model supports a wider range of characterizable models, making it a more realistic choice.
@@
Some studies have used graph samples to infer properties of the underlying manifold itself. These properties include the validity of the manifold assumption [58], the manifold dimension [59] and the complexity of these inferences [60, 61]. Other research has focused on prediction and classification using manifolds and manifold data, proposing various algorithms and methods. Impressive examples include the Isomap algorithm [62, 63, 64] and other manifold learning techniques [65]. These techniques aim to infer manifold properties without analyzing the generalization capabilities of GNNs operated on the sampled manifold.
@@
The transferability of GNNs has been extensively studied by examining the differences in GNN outputs across graphs of varying sizes as they converge to a limit model. This analysis, however, often lacks statistical generalization. Several studies have explored GNN transferability with graphon models, proving bounds on the differences in GNN outputs [15, 53, 54]. Other research has demonstrated how increasing graph size during GNN training can improve generalization to large-scale graphs [66]. The transferability of GNNs has also been investigated in the context of graphs generated from general topological spaces [67] and manifolds [35]. Furthermore, a novel graphop operator has been proposed as a limit model for both dense and sparse graphs, with proven transferability results [68]. Further research has focused on transfer learning for GNNs by measuring distances between graphs without assuming a limit model [69, 70]. Finally, a transferable graph transformer has been proposed and empirically validated [71].
@@ Appendix H Low Pass Filter Assumption
In the main results, we assume that the GNN and MNN are low-pass filters. This is a reasonable assumption because high-frequency signals on graphs or manifolds can fluctuate significantly between adjacent entries, leading to instability and learning challenges. We expect a degree of local homogeneity, which translates to low-frequency signals. This assumption is supported by empirical evidence in various domains, including opinion dynamics, econometrics, and graph signal processing [72, 73, 74].
Moreover, several other effective learning techniques, such as Principal Component Analysis (PCA) and Isomap, implicitly employ low-pass filtering. Therefore, we believe that the low-pass filter assumption is not restrictive and is well-supported by both practical applications and theoretical considerations.